OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 10

OpenStudy (anonymous):

@Turtledove

OpenStudy (uri):

4x+4=4x+10 4x-4x=10-4 0=6

OpenStudy (anonymous):

so its six

OpenStudy (anonymous):

so there is no solution

OpenStudy (uri):

There is no solution because 4x-4x got cancelled.

OpenStudy (lncognlto):

There really is no solution to this equation.

OpenStudy (texaschic101):

if your answer ends like this : 3 = 2, or 0 = 4...then their is no solution. if your answer ends like this : 2 = 2 or 5 = 5...then there is infinite solutions

OpenStudy (anonymous):

yes baby